OpenStudy (anonymous):

PLEASE HELP The product of two different perfect squares is 1,600.1,600 is also a perfect square. A) what is the square root of 1600? show your work..

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

ok il give a hint maybe it can gives clue :) \(\large 1600 = 16 \times 100\)

ganeshie8 (ganeshie8):

\(\large 1600 = 16 \times 100\) \(\large \sqrt{1600} = \sqrt{16 \times 100}\) \(\large = \sqrt{16} \times \sqrt{100}\)

ganeshie8 (ganeshie8):

can u take it from here ? :)

OpenStudy (anonymous):

no im still confused

ganeshie8 (ganeshie8):

\(\large 1600 = 16 \times 100\) \(\large \sqrt{1600} = \sqrt{16 \times 100}\) \(\large = \sqrt{16} \times \sqrt{100}\) \(\large = \sqrt{4^2} \times \sqrt{10^2}\)

ganeshie8 (ganeshie8):

now ?

ganeshie8 (ganeshie8):

\(\large 1600 = 16 \times 100\) \(\large \sqrt{1600} = \sqrt{16 \times 100}\) \(\large = \sqrt{16} \times \sqrt{100}\) \(\large = \sqrt{4^2} \times \sqrt{10^2}\) \(\large = 4 \times 10\) \(\large = 40\)

OpenStudy (anonymous):

ohh yeah

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

so my answer is 40 right ???

OpenStudy (anonymous):

ok thank you

ganeshie8 (ganeshie8):

np :)